POLARIZATION QUAMTUM BEAT SPECTROSCOPY OF NO_{2}: HYPERFINE LEVEL STRUCTURE IN THE YELLOW REGION(16000-19250 CM^{−1})

\underline{JU XIN} AND SCOTT A. REID, Department of Chemistry, Marquette University, Milwaukee, WI 53201-1881.

We report on the application of fluorescence-based polarization quantum beat spectroscopy to probe the excited hyperfine level structure of NO_{2} vibronic bands in the yellow region (16000-19250 cm^{−1}). Two general trends are observed when comparing our results with spectroscopic studies at lower energy. First, fluctuations in the Fermi-contact interaction dampen with increasing energy. Second, the magnitude of this interaction decreases with increasing enegy. We will dscuss possible explanations for and implication of these trends.